OpenStudy (anonymous):

How does a Tennis Player use Newtons three laws of motion? Medals will be given.

OpenStudy (anonymous):

@sammixboo @jane11509

OpenStudy (jane11509):

Yeah?

OpenStudy (jane11509):

I don't know o.o

OpenStudy (anonymous):

It's okay thank you though.

OpenStudy (anonymous):

By using projectile

OpenStudy (anonymous):

ΣF=ma

OpenStudy (anonymous):

In which we start off by saying Vx=Vocos(x) Vy=Vosin(x) and then start off by saying that a=-9.8m/s^2 cause the ball is in motion in space.

OpenStudy (anonymous):

Thats is ay ofcourse and ax=0 since ofcourse there is no acceleration along the x during its motion and then you start by integrating until you reach the displace x and y which will result to giving these 2 parametric equ X=Vo(t)cos(x) + xo Y=-1/2(g)t^2 + Vo(t)sin(x) +yo